problem:

The initial energy stored in the circuit in is zero. Find vy(t) for 120. 250 mH 80 mA 2002 0,00) +16 4F
Answer #1

tző 1 (0)=0A I volo=or at t=o Switch is closed - 250mH SomA Volt) = 16ME to be the onlot = 400 woahe Movie ) S60 Xawo It is uVo(t)= 16+ (A, cos(300t) + Assin (300t)) look 50 t=0 vo (0)=16+ A, = 0 R A =-16 dolos a icloty = 18,156 = 0 allsec dvoft? -40